A 98.0 g wooden block is initially at rest on a rough horizontal surface when a 12.2 g bullet is fired horizontally into (but does not go through) it. After the impact, the block–bullet combination slides 6.5 m before coming to rest. If the coefficient of kinetic friction between block and surface is 0.750, determine the speed of the bullet (in m/s) immediately before impact.
Answer #1

Solection : * We have M = 98 9 m = 12-29 x = 6-5 m el = 0.75 By kinematic equation - v2_u²=zas => 10)-(42) = 260.75*9.8) (6.5